OpenStudy (anonymous):

Find angle theta if cos = 4/7 and sin is negative...plz explain in detail I have no idea what I'm doing

OpenStudy (anonymous):

\[\cos \Theta = 4/7 \]\[\sin \Theta = negative\]

OpenStudy (anonymous):

i dont understand wat u r asking

OpenStudy (anonymous):

it's trigonometry i have to draw a graph of the angle